OpenStudy (anonymous):

the length of a rectangle is increasing at a rate of 8cm/s and its width is increasing at a rate of 3cm/s. When the length is 20cm and the width is 10cm, how fast is the area of the rectangle increasing? calculus question help pleasee

OpenStudy (shubhamsrg):

any attempt ?

OpenStudy (anonymous):

um yeah kind of

OpenStudy (shubhamsrg):

will you show us your attempt ?

OpenStudy (nurali):

dl/dt = 8 cm/s dw/dt = 3 cm/s A = lw dA/dt = l*dw/dt + w*dl/dt When l = 20 cm and w = 10 cm dA/dt = 20(3) + 10(8) dA/dt = 60 +80 dA/dt = ____________cm²/s